Are m1=7/4, m2=4/7 parallel, perpendicular or neither?
Answer by Alan3354(69443)

Are m1=7/4, m2=4/7 parallel, perpendicular or neither?
=====================
Parallel --- slopes are equal
Does m1 = m2?
-----
Perpendicular --- slopes are negative inverses, ie, the product of the 2 slopes = -1
Is m1*m2 = -1?